The length of a rectangle is 3 more than it's width. If the perimiter is 34, what is the length?
Answer by richard1234(7193)

Let the width be w and the length be 3+w. The perimeter is 34, so

2(w + (3+w)) = 34. The left-hand side simplifies to 2(3 + 2w) so

2(3 + 2w) = 34

3 + 2w = 17

w = 7, 3+w = 10 --> Length = 10

Checking, 2(7 + 10) = 34